The lengths of the sides of a right triangle are consecutive integers. Find the lengths of the three sides. (I know the answer is 3, 4, and 5 but how do I show the work?)
Answer by ewatrrr(24785)


Hi,
Let x represent the shortest side or smallest consecutive integers.
the other side (x+1) and the longest (the hypotenuse) (x + 2)
Appying the Pythagorean Theorem
x^2 + (x+1)^2 = (x+2)^2
x^2 + x^2 +2x + 1 = x^2 + 4x + 4
x^2 -2x -3 = 0
(x-3)(x+1) = 0
(x-3)=0
x = 3, the length of the shortest side
(x+1) = 0
x = -1 cannot use
3 , 4 , 5 are the length of the sides